LINESHAPE PARAMETERS OF THE OXYGEN A-BAND USING FREQUENCY-STABILIZED CAVITY RING-DOWN SPECTROSCOPY

\underline{D. J. ROBICHAUD} and M. OKUMURA, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena CA 91125; J. T. HODGES, National Institute of Standards and Technology, Gaithersburg, MD 20899; D. LISAK, Nicolaus Copernicus University, Torun, Poland; C. E. MILLER and L. R. BROWN, Jet Propulsion Laboratory, Pasadena, CA 91109.

Laboratory spectra of the oxygen A-band (b^{1}Σ_{g}^{+} ← X^{3}Σ_{g}^{−}) have been recorded using frequency-stabilized cavity ring-down spectroscopy (FS-CRDS) in the 12,900-13,080 cm^{−1} spectral region (20<N′′<40). High-resolution and high-sensitivity measurements of line shape parameters (intensities, pressure broadening and shifting) are reported. Implication of linemixing and collision-induced absorption to far-wing absorption will be discussed.